# Li Tana,b , Shengrong Wanga,c

### v1

@@ 2 Preliminaries
Let Assumption 2.4 hold, then, for any $x,y\in\mathbb{R}^{d}$, we have
$|b(x,y,\mu)|\leq$ $|b(x,y,\mu)-b(0,0,\mu)|+|b(0,0,\mu)|$
$\leq$ $K_{3}(1+|x|^{l}+|y|^{l})(|x|+|y|)+K_{5}(1+\mathbb{W}_{q}(\mu,\delta_{0}))$
$\leq$ $C(1+|x|^{l+1}+|y|^{l+1}+\mathbb{W}_{q}(\mu,\delta_{0})),$
where $C=K_{3}\vee K_{5}$. Moreover, by Young’s inequality, we get
$(x-D(y))b(x,y,\mu)=$ $(x-D(y)-0+D(0))(b(x,y,\mu)-b(0,0,\delta_{0})+b(0,0,\delta_{0}))$
$\leq$ $K_{2}(|x|^{2}+|y|^{2}+\mathbb{W}_{q}^{2}(\mu,\delta_{0}))+(|x|+D(y))|b(0,0,\delta_{0})|$
$\leq$ $C(1+|x|^{2}+|y|^{2}+\mathbb{W}_{q}^{2}(\mu,\delta_{0})),$
where $C=K_{2}\vee(K_{2}+1)\vee(K_{2}+\lambda)\vee(K_{2}+K_{5}^{2})$.
@@
[31, Lemma 4.1] Let $p>1$, $\varepsilon>0$ and $x,y\in\mathbb{R}$. Then
@@
Especially for $\varepsilon=((1-a)/a)^{p-1}$, where $p\geq 2$ and $10<a<$ 1, we get
@@ 3 Existence and Uniqueness
Since we are studying equations with super-linearly growing coefficients, it is necessary to consider the Carathéodory approximation. Now we define the Carathéodory approximation of (2.5) as follows. For every integer $n\geq 2/\tau$, define $X_{t}^{n}$ on $[-\tau,T]$ by
$X_{t}^{n}=\xi(t),\ \ t\in[-\tau,0],$
@@
(3.1) $X_{t}^{n}-D(X_{t-\tau}^{n})=\xi(0)-D(\xi(-\tau))+\int_{0}^{t}b\big{(}X_{s-1/n}^{n},X_{s-\tau}^{n},\mathcal{L}_{X_{s-1/n}^{n}}\big{)}{\rm d}s+\int_{0}^{t}\sigma\big{(}\mathcal{L}_{X_{s-1/n}^{n}}\big{)}{\rm d}B_{s}^{H},$
for $t\in[0,T]$. It is important to note that each $X_{t}^{n}$ can be determined explicitly by the stepwise iterated Itô integrals over the intervals $[0,1/n]$, $(1/n,2/n]$ etc. To get the final results, we give the following lemmas.
Lemma 3.1.
Let Assumptions 2.2-2.5 hold. Then for $\overline{p}>1/H$ and $n\geq 2/\tau$, we get
$\mathbb{E}\bigg{(}\sup_{0\leq t\leq T}|X_{t}^{n}|^{\overline{p}}\bigg{)}\leq C.$
Proof.
By Assumption 2.4 and Lemma 2.7, we get
$|X_{t}^{n}|^{\overline{p}}=$ $|D(X_{t-\tau}^{n})+X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}}$
$\leq$ $\frac{1}{\lambda^{p-1}}|D(X_{t-\tau}^{n})|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}|X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}}$
$\leq$ $\lambda|X_{t-\tau}^{n}|^{\overline{p}}+\frac{1}{(1-\lambda)^{p-1}}|X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}},$
for any $t\in[0,T]$. Then
$\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\leq$ $\lambda\parallel\xi\parallel^{p}+\lambda\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}\sup_{0\leq r\leq t}|X_{r}^{n}-D(X_{r-\tau}^{n})|^{\overline{p}},$
we may compute that
(3.2) $\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\leq\frac{\lambda}{1-\lambda}\parallel\xi\parallel^{\overline{p}}+\frac{1}{(1-\lambda)^{\overline{p}}}\sup_{0\leq r\leq t}|X_{r}^{n}-D(X_{r-\tau}^{n})|^{\overline{p}}$
By Assumption 2.3, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s}^{n}|^{\overline{p}}\bigg{)}$
$\leq$ $3^{\overline{p}-1}\mathbb{E}|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+3^{\overline{p}-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}b\left(X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\right){\rm d}r\bigg{\rvert}^{\overline{p}}\bigg{)}$
$+3^{\overline{p}-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\sigma\left(\mathcal{L}_{X_{r-1/n}^{n}}\right){\rm d}B_{r}^{H}\bigg{\rvert}^{\overline{p}}\bigg{)}$
$\leq$ $C+C\mathbb{E}\int_{0}^{t}\left[1+|X_{r-1/n}^{n}|^{(l+1)\overline{p}}+|X_{r-\tau}^{n}|^{(l+1){\overline{p}}}+\left(\mathbb{W}_{\theta}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\right)^{\overline{p}}\right]{\rm d}r$
$+C\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{\rvert}^{\overline{p}}\bigg{)}.$
Using the techniques of [12, Theorem 3.1]. Denote $\int_{s}^{t}(t-\beta)^{-\beta}(r-s)^{\beta-1}{\rm d}r:=C(\beta)$. Since for $\overline{p}H>1$, we can take some $\beta$ such that $1-H<\beta<1-1/{H}$, then by the stochastic Fubini theorem, the Hölder inequality, and combining with Lemma 2.4, we get
(3.3) $\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{\rvert}^{\overline{p}}\bigg{)}$
$\leq$ $\frac{C(\beta)^{-{\overline{p}}}}{({\overline{p}}-1-\beta{\overline{p}})^{{\overline{p}}-1}}t^{{\overline{p}}-1-\beta{\overline{p}}}\int_{0}^{t}\mathbb{E}\bigg{|}\int_{0}^{s}(s-r)^{\beta-1}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{|}^{{\overline{p}}}{\rm d}s$
$\leq$ $Ct^{{\overline{p}}-1-\beta{\overline{p}}}\int_{0}^{t}\left(\int_{0}^{s}(s-r)^{\frac{\beta-1}{H}}\parallel\sigma(\mathcal{L}_{X_{r-1/n}^{n}})\parallel^{\frac{1}{H}}{\rm d}r\right)^{\overline{p}H}{\rm d}s$
$\leq$ $Ct^{\overline{p}H-1}\int_{0}^{t}\parallel\sigma(\mathcal{L}_{X_{r-1/n}^{n}})\parallel^{\overline{p}}{\rm d}r$
$\leq$ $Ct^{\overline{p}H-1}L^{\overline{p}}\int_{0}^{t}\left[1+\mathbb{W}_{\theta}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\right]^{\overline{p}}{\rm d}r.$
Referring to the proof in [6, Proposition 3.4], for the $q$-Wasserstein metric, we get

### v2

@@ 2 Preliminaries
Let Assumptions 2.4-2.5 hold, then, for any $x,y\in\mathbb{R}^{d}$, we have
$|b(x,y,\mu)|\leq$ $|b(x,y,\mu)-b(0,0,\delta_{0})|+|b(0,0,\delta_{0})|$
$\leq$ $K_{3}[(1+|x|^{l}+|y|^{l})(|x|+|y|)+\mathbb{W}_{2}(\mu,\delta_{0})]+K_{5}$
$\leq$ $C(1+|x|^{l+1}+|y|^{l+1}+\mathbb{W}_{2}(\mu,\delta_{0})),$
where $C=2(K_{3}+K_{5})$. Moreover,
$\|\sigma(\mu)\|\leq$ $\|\sigma(\mu)-\sigma(\delta_{0})\|+\|\sigma(\delta_{0})\|\leq K_{4}\mathbb{W}_{2}(\mu,\delta_{0})+K_{5}\leq C(1+\mathbb{W}_{2}(\mu,\delta_{0})),$
where $C=K_{4}\vee K_{5}$. Further, let Assumptions 2.3-2.5 hold, then by the Young’s inequality, we get
$\langle x-D(y),b(x,y,\mu)\rangle=$ $\langle x-D(y)-0+D(0),b(x,y,\mu)-b(0,0,\delta_{0})+b(0,0,\delta_{0})\rangle$
$\leq$ $K_{2}(|x|^{2}+|y|^{2}+\mathbb{W}_{2}^{2}(\mu,\delta_{0}))+(|x|+|D(y)|)|b(0,0,\delta_{0})|$
$\leq$ $C(1+|x|^{2}+|y|^{2}+\mathbb{W}_{2}^{2}(\mu,\delta_{0})),$
where $C=(K_{2}+1/2)\vee K_{5}^{2}$.
For later use, we state the following lemmas without proof.
@@
[32, Lemma 4.1] Let $p>1$, $\varepsilon>0$ and $x,y\in\mathbb{R}$. Then
@@
Especially for $\varepsilon=[(1-a)/a]^{p-1}$ where $0<a<1$, we get
@@
Lemma 2.8.
[33, Lemma 7.1.2] Let $\beta>0$, $\gamma>0$ with $\beta+\gamma>1$. Suppose $a\geq 0$, $b\geq 0$, $u$ is nonnegative and $t^{\gamma-1}u(t)$ is locally integrable on $0\leq t\leq T$. If
$u(t)\leq a+b\int_{0}^{t}(t-s)^{\beta-1}s^{\gamma-1}u(s){\rm d}s,\ a.e.$
then
$u(t)\leq aE_{\beta,\gamma}((b\Gamma(\beta))^{1/v}t),$
where $v=\beta+\gamma-1>0$, and $E_{\beta,\gamma}(s)$ denotes the generalized Mittag-Leffler function defined by $E_{\beta,\gamma}(s)=\sum_{m=0}^{\infty}c_{m}s^{mv}$ with $c_{0}=1$, $\frac{c_{m+1}}{c_{m}}=\frac{\Gamma(mv+\gamma)}{\Gamma(mv+\gamma+\beta)}$ for all $m\geq 0$, here $\Gamma(\cdot)$ is the Gamma function.
Remark 2.9.
Considering the special situation of Lemma 2.8 with $\beta=2H(1/2<H<1)$, $\gamma=1$, $b=1$, it follows immediately that $u(t)$ is bounded, i.e., there exists a positive constant $C$ such that
$u(t)\leq C.$
Since for these parameters, we have $\beta+\gamma=2H+1>1$, so $v=2H>0$. The Mittag-Leffler function in Lemma 2.8 reduces to $E_{2H,1}(s)=\sum_{m=0}^{\infty}c_{m}s^{2Hm}$ where $s=[\Gamma(2H)]^{1/2H}t$. This series converges for all $t\in\mathbb{R}$, as shown by analyzing the asymptotic behavior of its coefficients. Using the Stirling’s approximation for the Gamma function, we derive
$\frac{c_{m+1}}{c_{m}}=\frac{\Gamma(2Hm+1)}{\Gamma(2Hm+1+2H)}\sim\frac{(2Hm)^{2Hm+1/2}e^{-2Hm}}{(2Hm+2H)^{2Hm+2H+1/2}e^{-(2Hm+2H)}}.$
Simplifying further, this ratio satisfies
$\frac{c_{m+1}}{c_{m}}\sim(2Hm)^{-2H}e^{2H}.$
Since for $2H>0$ we have
$\lim_{m\to\infty}(2Hm)^{-2H}e^{2H}=0,$
the ratio test guarantees that the series $\sum_{m=0}^{\infty}c_{m}s^{2Hm}$ has an infinite radius of convergence. Thus, $E_{2H,1}(\cdot)$ is bounded for all $t\in\mathbb{R}$, implying $u(t)\leq a\cdot E_{2H,1}(\cdot)\leq C.$
@@ 3 Existence and Uniqueness
Theorem 3.1.
Let Assumptions 2.2-2.5 hold. Then there exists a unique solution $\{X_{t}\}_{t\in[-\tau,T]}$ to (2.5). Moreover, for any $\overline{p}\geq 2$, we have
(3.1) $\mathbb{E}\bigg(\sup_{0\leq t\leq T}|X_{t}|^{\overline{p}}\bigg)\leq C,$
where $C$ is a positive constant depending on $T,H,\overline{p},\lambda,K_{i}$ and the initial data $\xi$.
Proof.
We note that the application of the fractional Itô formula in the following argument is justified by the Lipschitz conditions in Assumptions 2.3-2.5, which ensure the necessary regularity of the Picard iteration sequence $X_{t}^{(k)}$. Standard theory of Malliavin calculus for SDEs (e.g. [28]) guarantees that these iterates are $\phi$-differentiable and satisfy the integrability conditions required by Lemma 2.1. We refer to the proof procedure in [24] and divide the proof into 3 steps.
Step 1. Moment Boundness For $k\geq 0$ and $t\in[-\tau,0]$, define $X^{(k)}_{t}=\xi(t)$. For $k=0$ and $t\in[0,T]$, define $X^{(0)}_{t}\equiv\xi(0)$. For $k\geq 1$ and $t\in[0,T]$, $X_{t}^{(k)}$ is solved by
${\rm d}(X_{t}^{(k)}-D(X^{(k)}_{t-\tau}))=b(X_{t}^{(k)},X_{t-\tau}^{(k)},\mathcal{L}_{X_{t}^{(k-1)}}){\rm d}t+\sigma(\mathcal{L}_{X_{t}^{(k-1)}}){\rm d}B_{t}^{H},$
where $\mathcal{L}_{X_{t}^{(0)}}=\mathcal{L}_{\xi_{0}}$. For $i=1$, by Assumption 2.3 and Lemma 2.7, we get
$|X_{t}^{(1)}|^{\overline{p}}=$ $|D(X_{t-\tau}^{(1)})+X_{t}^{(1)}-D(X_{t-\tau}^{(1)})|^{\overline{p}}$
$\leq$ $\frac{1}{\lambda^{{\overline{p}}-1}}|D(X_{t-\tau}^{(1)})|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}|X_{t}^{(1)}-D(X_{t-\tau}^{(1)})|^{\overline{p}}$
$\leq$ $\lambda|X_{t-\tau}^{(1)}|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}|X_{t}^{(1)}-D(X_{t-\tau}^{(1)})|^{\overline{p}}.$
Then, we conclude that
(3.2) $\sup_{0\leq s\leq t}|X_{s}^{(1)}|^{\overline{p}}\leq\frac{\lambda}{1-\lambda}\|\xi\|^{\overline{p}}+\frac{1}{(1-\lambda)^{\overline{p}}}\sup_{0\leq s\leq t}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{\overline{p}}.$
Using the fractional Itô formula in Lemma 2.1, it follows that
$|X_{t}^{(1)}-D(X_{t-\tau}^{(1)})|^{\overline{p}}$
$\leq$ $|\xi(0)-D(\xi(-\tau))|^{\overline{p}}$
$+{\overline{p}}\int_{0}^{t}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{{\overline{p}}-2}\langle X_{s}^{(1)}-D(X_{s-\tau}^{(1)}),b(X_{s}^{(1)},X_{s-\tau}^{(1)},\mathcal{L}_{X_{s}^{(0)}})\rangle{\rm d}s$
$+\int_{0}^{t}{\overline{p}}({\overline{p}}-1)H(2H-1)|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{{\overline{p}}-2}\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\mathcal{L}_{X_{r}^{(0)}})\|^{2}{\rm d}r{\rm d}s$
$+{\overline{p}}\int_{0}^{t}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{{\overline{p}}-1}\|\sigma(\mathcal{L}_{X_{s}^{(0)}})\|{\rm d}B_{s}^{H}.$
$=$ $|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+Q_{1}^{(1)}(t)+Q_{2}^{(1)}(t)+Q_{3}^{(1)}(t).$
For any positive integer $N$, define the stopping time
$\rho_{N}^{(1)}=T\wedge\inf\{t\in[0,T]:|X_{t}^{(1)}|\geq N\}.$
It yields that $\rho_{N}^{(1)}\to T$ a.s. as $N\to\infty$. Applying Assumptions 2.3 and 2.4, the Young’s inequality, we see
$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}Q_{1}^{(1)}(s)\bigg)$
$=$ ${\overline{p}}\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{{\overline{p}}-2}\langle X_{s}^{(1)}-D(X_{s-\tau}^{(1)}),b(X_{s}^{(1)},X_{s-\tau}^{(1)},\mathcal{L}_{X_{s}^{(0)}})\rangle{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}\bigg[1+|X_{s}^{(1)}|^{2}+|X_{s-\tau}^{(1)}|^{2}+\mathbb{W}_{2}^{2}(\mathcal{L}_{X_{s}^{(0)}},\delta_{0})\bigg]^{{\overline{p}}/2}{\rm d}s$
$\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}|^{\overline{p}}\bigg){\rm d}r.$
Since Assumption 2.5 implies $\|\sigma(\mathcal{L}_{X_{r}^{(0)}})\|\leq C\big(1+\mathbb{W}_{2}(\mathcal{L}_{X_{r}^{(0)}},\delta_{0})\big)\leq C$, this together with the Young’s inequality, the Jensen inequality, and the Fubini theorem, it results in
$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}Q_{2}^{(1)}(s)\bigg)$
$=$ $\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}{\overline{p}}({\overline{p}}-1)H(2H-1)|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{{\overline{p}}-2}\int_{0}^{s}(s-r)^{2H-2}\|\sigma(\mathcal{L}_{X_{r}^{(0)}})\|^{2}{\rm d}r{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}\int_{0}^{s}(s-r)^{2H-2}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{{\overline{p}}-2}(1+\mathbb{W}_{2}(\mathcal{L}_{X_{r}^{(0)}},\delta_{0}))^{2}{\rm d}r{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{\overline{p}}\bigg(\int_{0}^{s}(s-r)^{2H-2}{\rm d}r\bigg){\rm d}s+C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}\int_{0}^{s}(s-r)^{2H-2}{\rm d}r{\rm d}s$
$=$ $C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{\overline{p}}\bigg(\int_{0}^{s}(s-r)^{2H-2}{\rm d}r\bigg){\rm d}s+C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}\int_{r}^{t}(s-r)^{2H-2}{\rm d}s{\rm d}r$
$=$ $C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}\frac{s^{2H-1}}{2H-1}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}\frac{(t-r)^{2H-1}}{2H-1}{\rm d}r.$
Combining with Assumption 2.3 and the fact that $1/2<H<1$, we obtain
$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}Q_{2}^{(1)}(s)\bigg)\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}|^{\overline{p}}\bigg){\rm d}r+C\mathbb{E}\int_{0}^{t\wedge\rho_{N}^{(1)}}(t-r)^{2H-1}{\rm d}r.$
For $Q_{3}^{(1)}(s)$, we use the fractional Burkholder-Davis-Gundy (BDG) inequality. For $H\in(1/2,1)$ and $p\geq 1$, there exists a constant $C_{H,p}>0$ such that
(3.3) $\mathbb{E}\left[\sup_{0\leq s\leq t}\left|\int_{0}^{s}u_{r}{\rm d}B_{r}^{H}\right|^{p}\right]\leq C_{H,p},\mathbb{E}\left[\left(\int_{0}^{t}|u_{r}|^{1/H}{\rm d}r\right)^{pH}\right].$
Applying this with $p=1$ and $u_{r}=|X_{r}^{(1)}-D(X_{r-\tau}^{(1)})|^{{\overline{p}}-1}\|\sigma(\mathcal{L}_{{X_{r}^{(0)}}})\|$, and noticing the fact that Assumption 2.5 yields $\|\sigma(\mathcal{L}_{X_{r}^{(0)}})\|\leq C\big(1+\mathbb{W}_{2}(\mathcal{L}_{X_{r}^{(0)}},\delta_{0})\big)\leq C$, we get
$\mathbb{E}\left(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}\left|Q_{3}^{(1)}(s)\right|\right)=\overline{p}\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}\left|\int_{0}^{s}u_{r}{\rm d}B_{r}^{H}\right|\bigg)\leq C\mathbb{E}\left[\left(\int_{0}^{t\wedge\rho_{N}^{(1)}}|u_{r}|^{1/H}{\rm d}r\right)^{H}\right]$
$\leq$ $C\mathbb{E}\left[\left(\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{r}^{(1)}-D(X_{r-\tau}^{(1)})|^{({\overline{p}}-1)/H}{\rm d}r\right)^{H}\right].$
For any $0<\epsilon<1$, the Young’s inequality with exponents $\frac{\overline{p}}{\overline{p}-1}$ and its conjugate gives
$|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{\overline{p}-1}\leq\epsilon|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{\overline{p}}+C_{\epsilon},$
where $C_{\epsilon}$ depends on $\epsilon,\overline{p},H$. Integrating over $0,t\wedge\rho_{N}^{(1)}]$, then
$\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{(\overline{p}-1)/H}\mathrm{d}r\leq 2^{1/H-1}\epsilon^{1/H}\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{\overline{p}/H}\mathrm{d}r+2^{1/H-1}C_{\epsilon}^{1/H}t.$
Since $H\in(1/2,1)$, we use the elementary inequality to get
$\left(\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{(\overline{p}-1)/H}\mathrm{d}r\right)^{H}\leq$ $2^{1-H}\epsilon\left(\int_{0}^{t\wedge\rho_{N}^{(1)}}|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{\overline{p}/H}\mathrm{d}r\right)^{H}+2^{1-H}C_{\epsilon}t^{H}$
$\leq$ $2^{1-H}\epsilon t^{H}\cdot\sup_{0\leq r\leq t\wedge\rho_{N}^{(1)}}|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{\overline{p}}+2^{1-H}C_{\epsilon}t^{H}.$
Taking expectations and using Assumption 2.3
$\mathbb{E}\left(\sup_{0\leq r\leq t\wedge\rho_{N}^{(1)}}|X_{r}^{(1)}-D\big(X_{r-\tau}^{(1)}\big)|^{\overline{p}}\right)\leq C\left(1+\mathbb{E}\left(\sup_{0\leq r\leq t\wedge\rho_{N}^{(1)}}\left|X_{r}^{(1)}\right|^{\overline{p}}\right)\right).$
Combining all estimations, we arrive at
(3.4) $\mathbb{E}\left(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}\left|Q_{3}^{(1)}(s)\right|\right)\leq C\epsilon\mathbb{E}\left(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}\left|X_{s}^{(1)}\right|^{\overline{p}}\right)+C_{\epsilon}.$
Choosing $\epsilon$ sufficiently small so that $C\epsilon<\frac{1}{2}$, and combining with the estimates for $Q_{1}^{(1)}(t)$ and $Q_{2}^{(1)}(t)$, we get
$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}-D(X_{s-\tau}^{(1)})|^{\overline{p}}\bigg)\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}|^{\overline{p}}\bigg){\rm d}r.$
By the Gronwall inequality together with (3.2), we derive
$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(1)}}|X_{s}^{(1)}|^{\overline{p}}\bigg)\leq$ $C.$
The Fatou lemma then leads to
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{(1)}|^{\overline{p}}\bigg)\leq C.$
For $i=2$, we estimate the corresponding $Q_{1}^{(2)}(s),Q_{2}^{(2)}(s)$ and $Q_{3}^{(2)}(s)$ as follows
$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(2)}}Q_{1}^{(2)}(s)\bigg)\leq$ $C\bigg[\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(2)}}|X_{s}^{(2)}|^{\overline{p}}\bigg){\rm d}r+\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(2)}}|X_{s}^{(1)}|^{\overline{p}}\bigg){\rm d}r+1\bigg],$
$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(2)}}Q_{2}^{(2)}(s)\bigg)\leq$ $C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(2)}}|X_{s}^{(2)}|^{\overline{p}}\bigg){\rm d}r+C\int_{0}^{t}(t-r)^{2H-1}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(2)}}|X_{s}^{(1)}|^{\overline{p}}\bigg){\rm d}r,$
@@
$\mathbb{E}\bigg(\sup_{0\leq s\leq t\wedge\rho_{N}^{(2)}}Q_{3}^{(2)}(s)\bigg)\leq$ $C\epsilon\mathbb{E}\left(\sup_{0\leq s\leq t\wedge\rho_{N}^{(2)}}\left|X_{s}^{(2)}\right|^{\overline{p}}\right)+C_{\epsilon}\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(2)}}|X_{s}^{(2)}|^{\overline{p}}\bigg){\rm d}r$
$+C\int_{0}^{t}\mathbb{E}\bigg(\sup_{0\leq s\leq r\wedge\rho_{N}^{(2)}}|X_{s}^{(1)}|^{\overline{p}}\bigg){\rm d}r.$
Consequently, choosing $\epsilon$ sufficiently small so that $C\epsilon<\frac{1}{2}$ and using the Fatou lemma, the Gronwall inequality, it holds
$\mathbb{E}\bigg(\sup_{0\leq s\leq t}|X_{s}^{(2)}|^{\overline{p}}\bigg)\leq C.$
Finally, (3.1) can be shown by induction.
Step 2. Existence Define the difference processes $\Delta_{s}^{(k+1),(k)}=X^{(k+1)}_{s}-X_{s}^{(k)}$, $\Upsilon_{s}^{(k+1),(k)}=X^{(k+1)}_{s}-D(X_{s-\tau}^{(k+1)})-X_{s}^{(k)}+D(X_{s-\tau}^{(k)})$. By Lemma 2.7, we have
$|\Delta_{s}^{(k+1),(k)}|^{\overline{p}}\leq\lambda|\Delta_{s-\tau}^{(k+1),(k)}|^{\overline{p}}+\frac{1}{(1-\lambda)^{\overline{p}-1}}|\Upsilon_{s}^{(k+1),(k)}|^{\overline{p}},$
Taking supremum and expectation gives